# Noise-Robustness Through Noise: Asymmetric LoRA Adaption with Poisoning Expert

### v3

@@ 1 Introduction
Recently, pre-trained language models (PrLMs) have demonstrated remarkable success across various natural language processing tasks. To further enhance model performance on downstream tasks, researchers typically employ domain-specific corpora for targeted fine-tuning of pre-trained models. However, this mainstream solution critically depends on the quality of training data in both pre-training and fine-tuning stages, though with distinct emphases. Different from the pre-training stage, where models acquire general linguistic knowledge through exposure to large-scale corpora, the fine-tuning stage requires models to adapt to specific downstream tasks using comparatively smaller datasets, thereby imposing stricter requirements on data quality.
In downstream NLP applications, noise poses multiple critical challenges: 1) labeling errors, syntactic irregularities, and extraneous content disrupt the model’s ability to capture and learn effective features, leading to unstable training and hindering the establishment of precise, robust knowledge systems Mao et al. (2023); Xu et al. (2024); 2) noise weakens model generalization when encountering unseen data, impairing flexible knowledge transfer Ahn et al. (2024); 3) noisy data may introduce biases, causing models to disproportionately emphasize or neglect specific categories, compromising fairness and interpretability during fine-tuning stage Zhao et al. (2024). These issues collectively manifest as drastic performance degradation in downstream tasks.
@@
Extensive experiments on benchmark datasets across mainstream NLP tasks demonstrate that our method achieves strong performance without relying on processed clean data, addressing limitations in existing denoising methods. A subsequent series of experiments further validate the fitness between our asymmetric LoRA architecture and noise-robust enhancement. Notably, since LoPE is based on a low-rank parameter-efficient fine-tuning framework, it maintains noise robustness while also achieving high efficiency. The main contributions of this paper are summarized as follows:
@@ 2.2 Development of LoRA
LLMs have attracted considerable attention due to their powerful natural language processing capabilities Kenton and Toutanova (2019). However, the massive parameters of these models make full-parameter tuning prohibitively expensive. Low-Rank Adaptation (LoRA) Hu et al. (2022) decomposes weight updates into trainable low-rank matrices ($A$, $B$), maintaining performance while drastically reducing parameters and computation, establishing it as a key parameter-efficient fine-tuning (PEFT) Li and Liang (2021) technique. Recent innovations integrate LoRA with Mixture-of-Experts (MoE) architectures. MoE-LoRA Dou et al. (2024) synthesizes complementary strengths: MoE’s specialized expert networks with gating mechanisms synergize with LoRA’s low-rank adaptations, enabling cross-domain adaptability. This hybrid approach mitigates multitask interference while preserving parameter efficiency through task-specific adaptor coordination. Li et al. (2024) integrated LoRA experts into the feedforward network (FFN) layers of a frozen pre-trained model, improving task adaptability while addressing imbalance issues in MoE models. To address the performance degradation of LoRA in complex corpora while maintaining parameter efficiency, Tian et al. (2024) developed HydraLoRA, an asymmetric structure that partitions LoRA modules into specialized experts with MoE-based routing, improving efficiency and adaptability in heterogeneous corpora. The asymmetric LoRA structure’s unique design enables explicit noise modeling and separates noise-processing modules. This provides a foundation framework for achieving noise robustness through the generated noise.
@@ 3 Methodology
Figure 2: The LoPE pipeline consists of two stages: fine-tuning and inference stages. Fine-tuning Stage I: HyNoIse is to enhance the poisoning expert’s noise understanding. Fine-tuning Stage II: Freezing the poisoning expert while fine-tuning the remaining experts and shared matrix. Inference stage: Masking the poisoning expert with noise-affect knowledge, allowing the remaining experts to generate output with relatively clean knowledge.
@@ 3.1 Asymmetric LoRA architecture
Here we first introduce the asymmetric LoRA architecture Tian et al. (2024) we employed. This architecture utilizes a centrally shared matrix ${A}$ and multiple distinct matrices ${B}_{i}$, shown as follows: $\textit{W}=W_{0}+\sum_{i=1}^{N}\omega_{i}B_{i}A$. Compared to symmetric architectures, asymmetric configurations demonstrate superior advantages in mitigating parameter redundancy and enhancing computational efficiency. By sharing a central matrix ${A}$, the architecture significantly reduces learnable parameters while maintaining representational capacity, yielding a more lightweight architecture. The implementation of multiple matrices ${B}_{i}$ enables flexible assignment of specialized adaptive modules for distinct objectives.
Because of these relatively independent experts, our approach, which involves isolating a dedicated module within the model to separate and process noise, becomes technically feasible. In LoPE, matrix $B$ is treated as the expert adapter. Furthermore, a randomly selected matrix $B$ is designated as a dedicated poisoning expert, denoted as $B_{D}$. This expert shares the same structure as the other experts but undergoes a targeted modification of its capabilities through the process described in the following sections, with the aim of identifying and processing noise. Its forward propagation is as follows: $\textit{W}=W_{0}+(\omega_{D}B_{D}+\sum_{i=1}^{N-1}\omega_{i}B_{i})A$.
@@ 3.2.3 Fine-tuning Stage II: Dynamically Compensated Expert Synergy
[…]
@@ 4.3 Main Results
[…]
@@
LoPE(r=4) ${\textbf{44.42}}_{\pm 0.18}$ ${\textbf{76.28}}_{\pm 0.38}$ ${\textbf{49.03}}_{\pm 0.41}$ ${\textbf{13.72}}_{\pm 0.34}$ ${\textbf{60.49}}_{\pm 0.27}$ ${\textbf{35.45}}_{\pm 0.33}$ ${\textbf{52.88}}_{\pm 0.20}$ ${\textbf{56.84}}_{\pm 0.46}$ 0.062 1 3
In this section, we evaluate LoPE’s effectiveness through extensive experiments and compare it with several mainstream PEFT methods on LLaMA2-7b Touvron et al. (2023). Table 1 presents the performance of LoPE alongside the baseline methods. MMLU represents the average accuracy across all tasks. NSET, SBH, and History represent the average accuracy of tasks within each respective set. The specific accuracy for each sub-task can be found in Appendix D. We first compared LoPE with HydraLoRA on the $\boldsymbol{Orig}$ datasets mentioned in Section 4.2. In the original environment, LoPE exhibited only a limited improvement across these datasets. Given that the $\boldsymbol{Orig}$ datasets are nearly noise-free with minimal inherent noise, LoPE’s noise adaptation capability was not significantly demonstrated. Subsequently, we conducted experiments on the $\boldsymbol{Nois}$ datasets. The results indicate that LoPE exhibits robustness and adaptability across most tasks. Notably, LoPE excelled across three QA tasks, achieving a 4.23% improvement on the ARC-e. Consistent with our theoretical analysis, LoPE also showed a 1.89% improvement on the GSM8K task. Moreover, LoPE outperformed baseline methods in the MMLU benchmark (average accuracy of 57 sub-tasks), achieving a 1.34% average improvement over HydraLoRA. Specifically, as the sub-task of MMLU, NSET, similar to GSM8K, often involves precise computations, LoPE maintained coherent reasoning under noise, whereas traditional methods struggled. Furthermore, LoPE also outperforms other methods in complex tasks such as SBH and History, which involve long texts and logical reasoning. In conclusion, LoPE overcomes the limitations of conventional denoising methods, achieving enhanced robustness through noise injection. Notably, our method maintains the same time complexity of $O(n^{2})$ as other fine-tuning methods, ensuring high computational efficiency. In Table 1, we also explore LoPE’s sensitivity to parameter configurations by examining the impact of varying the rank and the number of experts, as well as comparing its parameter scale with other methods.
@@ 5.1 Does Higher HyNoise Ratio Enhance Performance?
In Section 4.2, while our baseline experiments employed a 5% HyNoIse ratio, we investigated whether an increased noise ratio during Stage I enhances noise robustness. We evaluated four HyNoIse ratios (3.5%, 5%, 10%, 15%), with comparative results visualized in Figure 4. Contrary to potential expectations, our analysis reveals a non-monotonic relationship between HyNoIse intensity and model robustness. Excessive noise injection (i.e. $\geq$ 10%) degrades the router’s discriminative capability for noise patterns, as evidenced by reduced expert specialization. We consider that this was due to overwhelming data distortion, which prevents effective differentiation between clean and noisy inputs, ultimately compromising expert weight allocation.
@@ 5.2 Can the Poisoning Expert Truly Accomplish Its Task?
Method PE=3,NE=1 PE=3,NE=2 PE=2,NE=2 Mask(PE=3,NE=1) Not Mask(PE=3,NE=1)
@@ 6 Conclusion
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A Dataset
A.1 Fine-tuning Dataset
A.2 Evaluation Dataset
Table 5: Task categories and corresponding tasks
Task Category Tasks
Appendix B Baseline, Configurations and Backbone LLMs
B.1 Baseline
B.2 Experiment Configurations
B.3 Backbone LLMs
Appendix C Limitations
Appendix D Specific accuracy for sub-task
Task Model
NSET LoPE HydraLoRA
College Biology 44.44 43.75
College Chemistry 32.00 30.00
College Computer Science 37.00 37.00
College Mathematics 38.00 34.00
College Physics 16.67 18.63
Electrical Engineering 42.76 45.52
Astronomy 40.79 42.76
Abstract Algebra 32.00 30.00
Machine Learning 38.39 36.61
Computer Security 58.00 58.00
High School Chemistry 35.96 32.51
Elementary Mathematics 29.89 27.51
High School Physics 28.48 29.14
High School Computer Science 37.00 37.00
High School Statistics 30.56 25.00
High School Biology 47.42 49.35
High School Mathematics 29.26 28.89
Conceptual Physics 44.26 42.13
Task Model
SBH LoPE HydraLoRA
Moral Scenarios 27.71 24.69
Professional Psychology 43.14 43.63
World Religions 66.67 66.08
Philosophy 59.49 58.84
Moral Disputes 49.13 51.16
High School Government and Politics 66.84 68.39
Sociology 61.19 64.18
High School Psychology 61.47 61.28
Task Model
History LoPE HydraLoRA
High European History 60.61 63.64
Prehistory 47.22 47.22
High School World History 62.45 59.49
High School US History 56.37 53.43
Task Model
NSET LoPE HydraLoRA
College Biology 45.14 43.75
College Chemistry 31.00 29.00
College Computer Science 38.00 33.00
College Mathematics 33.00 40.00
College Physics 17.65 15.69
Electrical Engineering 38.62 33.79
Astronomy 42.11 38.16
Abstract Algebra 32.00 34.00
Machine Learning 38.39 37.50
Computer Security 53.00 52.00
High School Chemistry 33.50 33.00
Elementary Mathematics 26.72 27.51
High School Physics 31.13 27.81
High School Computer Science 41.00 38.00
High School Statistics 21.30 23.61
High School Biology 46.13 47.10
High School Mathematics 28.52 29.63
Conceptual Physics 40.85 41.70
Task Model
SBH LoPE HydraLoRA
Moral Scenarios 26.03 25.92
Professional Psychology 44.28 44.12
World Religions 63.74 64.33
Philosophy 56.27 56.91
Moral Disputes 48.84 48.55
High School Government and Politics 65.28 64.25
Sociology 58.21 56.22
High School Psychology 60.37 58.53
Task Model
History LoPE HydraLoRA
High European History 63.61 61.21
Prehistory 49.38 47.53
High School World History 59.92 59.07
High School US History 54.41 55.39

### v5

# Noise-Robustness Through Noise: A Framework combining Asymmetric LoRA with Poisoning MoE

@@ 1 Introduction
Recently, pre-trained language models (PrLMs) have demonstrated remarkable success across various natural language processing tasks Gao et al. (2025); Bianchi et al. ; Guo et al. (2025); Yang et al. . To further enhance model performance on downstream tasks, researchers typically employ domain-specific corpora for targeted fine-tuning of pre-trained models Li et al. (2024a). However, this mainstream solution critically depends on the quality of training data in both pre-training and fine-tuning stages, though with distinct emphases. Different from the pre-training stage, where models acquire general linguistic knowledge through exposure to large-scale corpora, the fine-tuning stage requires models to adapt to specific downstream tasks using comparatively smaller datasets, thereby imposing stricter requirements on data quality Mai et al. (2024); Shi and Lipani (2023).
In downstream NLP applications Ou et al. (2025), noise poses multiple critical challenges: 1) labeling errors, syntactic irregularities, and extraneous content disrupt the model’s ability to capture and learn effective features, leading to unstable training and hindering the establishment of precise, robust knowledge systems Mao et al. (2023); Xu et al. (2024); 2) noise weakens model generalization when encountering unseen data, impairing flexible knowledge transfer Ahn et al. (2024); 3) noisy data may introduce biases, causing models to disproportionately emphasize or neglect specific categories, compromising fairness and interpretability during fine-tuning stage Zhao et al. (2024). These issues collectively manifest as drastic performance degradation in downstream tasks.
@@
Extensive experiments on benchmark datasets across mainstream NLP tasks demonstrate that our method achieves strong performance without relying on processed clean data, addressing limitations in existing denoising methods. A subsequent series of experiments further validates the fitness between our asymmetric LoRA architecture and noise-robust enhancement. Notably, since LoPE is based on a low-rank parameter-efficient fine-tuning framework, it maintains noise robustness while also achieving high efficiency. The main contributions of this paper are summarized as follows:
@@ 2.2 Development of LoRA
LLMs have attracted considerable attention due to their powerful natural language processing capabilities Kenton and Toutanova (2019). However, the massive parameters of these models make full-parameter tuning prohibitively expensive. Low-Rank Adaptation (LoRA) Hu et al. (2022) decomposes weight updates into trainable low-rank matrices ($A$, $B$), maintaining performance while drastically reducing parameters and computation, establishing it as a key parameter-efficient fine-tuning (PEFT) Li and Liang (2021) technique. Recent innovations integrate LoRA with Mixture-of-Experts (MoE) architectures. MoE-LoRA Dou et al. (2024) synthesizes complementary strengths: MoE’s specialized expert networks with gating mechanisms synergize with LoRA’s low-rank adaptations, enabling cross-domain adaptability. This hybrid approach mitigates multitask interference while preserving parameter efficiency through task-specific adaptor coordination. Li et al. (2024b) integrated LoRA experts into the feedforward network (FFN) layers of a frozen pre-trained model, improving task adaptability while addressing imbalance issues in MoE models. To address the performance degradation of LoRA in complex corpora while maintaining parameter efficiency, Tian et al. (2024) developed HydraLoRA, an asymmetric structure that partitions LoRA modules into specialized experts with MoE-based routing, improving efficiency and adaptability in heterogeneous corpora. The asymmetric LoRA structure’s unique design enables explicit noise modeling and separates noise-processing modules. This provides a foundation framework for achieving noise robustness through the generated noise.
@@ 3 Methodology
Figure 2: The LoPE pipeline consists of two stages: fine-tuning and inference stages. Fine-tuning Stage I: HyNoIse is to enhance the poisoning expert’s noise understanding. Fine-tuning Stage II: Freezing the poisoning expert while fine-tuning the remaining experts and the shared matrix. Inference stage: Masking the poisoning expert with noise-affect knowledge, allowing the remaining experts to generate outputs that are relatively robust to noise.
@@ 3.1 Asymmetric LoRA architecture
Here we first introduce the asymmetric LoRA architecture Tian et al. (2024) we employed. This architecture utilizes a centrally shared matrix A{A} and multiple distinct matrices ${B}_{i}$, shown as follows: $\textit{W}=W_{0}+\sum_{i=1}^{N}\omega_{i}B_{i}A$. Compared to symmetric architectures, asymmetric configurations demonstrate superior advantages in mitigating parameter redundancy and enhancing computational efficiency. By sharing a central matrix A{A}, the architecture significantly reduces learnable parameters while maintaining representational capacity, yielding a more lightweight architecture. The implementation of multiple matrices ${B}_{i}$ enables flexible assignment of specialized adaptive modules for distinct objectives.
Because of these relatively independent experts, our approach, which involves isolating a dedicated module within the model to separate and process noise, becomes technically feasible. In LoPE, the matrix $B$ is treated as the expert adapter. Furthermore, a randomly selected matrix $B$ is designated as a dedicated poisoning expert, denoted as $B_{D}$. This expert shares the same structure as the other experts but undergoes a targeted modification of its capabilities through the process described in the following sections, with the aim of identifying and processing noise. Its forward propagation is as follows: $\textit{W}=W_{0}+(\omega_{D}B_{D}+\sum_{i=1}^{N-1}\omega_{i}B_{i})A$.
@@ 3.2.3 Fine-tuning Stage II: Dynamically Compensated Expert Synergy
We introduce a router network (gating function) to dynamically allocate expert adapters’ contributions. It processes the token representation $x$ through a fully connected layer and softmax to compute expert weights $\textit{$\omega_{i}$}=\textit{softmax}({W_{gate}}^{T}x)$, where $W_{gate}\in\mathbb{R}^{r\times N}$ denotes the router’s parameters. To address the interdependency disruption caused by expert masking in the inference stage, we propose the Dynamically Compensated Expert Synergy (DyCompEnSate) mechanism. Our analysis reveals that joint expert training establishes learned inter-expert correlations, where the direct mask of ${B}_{D}$ during inference has the potential to disrupt these dependencies. DyCompEnSate dynamically preserves these relationships through dependency-aware weight compensation. Formally, we construct an expert dependency matrix ${\Theta}\in\mathbb{R}^{N\times N}$, where the element $\theta_{ij}$ quantifies the similarity of output between experts $i$ and $j$. The poisoning expert dependencies $\theta_{iD}$ are computed via output similarity analysis: $\theta_{iD}=\textit{sim}({y}_{i},{y}_{D}),\quad i=1,2,\ldots,N,i\neq D$, where $\textit{sim}(\cdot,\cdot)$ denotes cosine similarity metric. These dependencies compensate expert contributions through dynamic weight amplification:
@@ 4.3 Main Results
Table 1: The comparative performance (%) of different PEFT methods across various datasets. The methods marked with a $\dagger$} are fine-tuned with the $\boldsymbol{Nois}$ datasets, while all subsequent methods are fine-tuned based on $\boldsymbol{Orig}$ datasets. r denotes the rank, #A represents the number of shared matrices, #B represents the number of $B$ matrices involved in the inference stage.
@@
LoPE(r=4)† ${44.42}_{\pm 0.18}$ ${76.28}_{\pm 0.38}$ ${49.03}_{\pm 0.41}$ ${13.72}_{\pm 0.34}$ ${60.49}_{\pm 0.27}$ ${35.45}_{\pm 0.33}$ ${52.88}_{\pm 0.20}$ ${56.84}_{\pm 0.46}$ 0.062 1 3
LoPE(r=8) ${\textbf{44.82}}_{\pm 0.24}$ ${\textbf{76.83}}_{\pm 0.35}$ ${\textbf{49.90}}_{\pm 0.27}$ ${\textbf{14.31}}_{\pm 0.38}$ ${\textbf{62.36}}_{\pm 0.47}$ ${\textbf{35.79}}_{\pm 0.13}$ ${\textbf{53.71}}_{\pm 0.32}$ ${\textbf{58.02}}_{\pm 0.34}$ 0.124 1 3
In this section, we evaluate LoPE’s effectiveness through extensive experiments and compare it with several mainstream PEFT methods on LLaMA2-7b Touvron et al. (2023). Table 1 presents the performance of LoPE alongside the baseline methods. MMLU represents the average accuracy across all tasks. NSET, SBH, and History represent the average accuracy of tasks within each respective set. The specific accuracy for each sub-task can be found in Appendix D.2. We first compared LoPE with HydraLoRA on the $\boldsymbol{Orig}$ datasets mentioned in Section 4.2. In the original environment, LoPE exhibited only a limited improvement across these datasets. Given that the $\boldsymbol{Orig}$ datasets are nearly noise-free with minimal inherent noise, LoPE’s noise adaptation capability was not significantly demonstrated. Subsequently, we conducted experiments on the $\boldsymbol{Nois}$ datasets. The results indicate that LoPE exhibits robustness and adaptability across most tasks. Notably, LoPE excelled across three QA tasks, achieving a 4.89% improvement on the ARC-e with a rank of 8. Consistent with our theoretical analysis, LoPE also showed an improvement on the GSM8K task. Moreover, LoPE outperformed baseline methods in the MMLU benchmark (average accuracy of 57 sub-tasks), achieving a 1.34% average improvement over HydraLoRA with a rank of 4. Specifically, as the sub-task of MMLU, NSET, similar to GSM8K, often involves precise computations, LoPE maintained coherent reasoning under noise, whereas traditional methods struggled. Furthermore, LoPE also outperforms other methods in complex tasks such as SBH and History, which involve long texts and logical reasoning. In conclusion, LoPE overcomes the limitations of conventional denoising methods, achieving enhanced robustness through noise injection. Notably, our method maintains the same time complexity of $O(n^{2})$ as other fine-tuning methods, ensuring high computational efficiency. In Table 1, we also explore LoPE’s sensitivity to parameter configurations by examining the impact of varying the rank and the number of experts, as well as comparing its parameter scale with other methods. Higher ranks generally lead to better performance; However, increasing the rank also incurs higher parameter and computational costs, thus requiring a balance between scalability and effectiveness. In the following experiments, the rank is set to 4 by default, with three normal experts and one poisoning expert.
@@ 4.4.1 Ablations on HyNoIse
To further investigate whether LoPE’s strong performance depends on the consistency between the noise types in the $\boldsymbol{Nois}$ dataset and the discrete noise in HyNoIse, we conducted additional experiments introducing inconsistent noise injection between the dataset construction and the discrete noise in HyNoIse. The results in Table 3 show that when different operations are applied for LoPE training and dataset construction, all LoPE variants achieve considerable improvements over the baseline HydraLoRA. This demonstrates that LoPE’s effectiveness generalizes across different noise categories and does not strongly depend on noise type consistency.
Moreover, the observed performance gains on the $\boldsymbol{Orig}$ dataset (Table 1) further indicate that LoPE is effective and generalizable to naturally occurring noisy data. It is worth noting that HyNoIse also injects continuous noise, which contributes to enhancing the model’s robustness under multiple types of noise.
Table 3: Performance comparison of LoPE under inconsistent noise injection and HydraLoRA. Each row represents a distinct noise operation. Abbreviations used: NCI (Noise Character Insertion), WOS (Word Order Shuffling), CD (Character Deletion), and WR (Word Replacement), representing four types of noise operations.
Model Dataset HyNoIse MMLU PIQA SIQA GSM8K
LoPE Orig + NCI Continuous + WOS 44.58 75.15 50.11 14.92
HydraLoRA Orig + NCI None 44.24 75.03 49.75 12.78
LoPE Orig + WOS Continuous + CD 45.09 75.84 50.55 14.64
HydraLoRA Orig + WOS None 44.37 75.21 49.65 12.34
LoPE Orig + WR Continuous + NCI 45.13 75.35 50.17 14.73
HydraLoRA Orig + WR None 44.31 75.07 49.73 12.39
We further applied HyNoIse directly to the original HydraLoRA framework (Table 4) and observed a notable performance drop. This is mainly due to the absence of the division of labor between normal experts and poisoning experts, and the multi-stage differentiated training, causing noisy and clean data to equally affect the model from the start. Therefore, LoPE’s key contribution is its architecture and multi-stage process, enabling explicit noise filtering and robustness.
Table 4: Performance comparison of LoPE and HydraLoRA (with/without HyNoIse) across benchmarks. The methods marked with a $\dagger$} are fine-tuned with the $\boldsymbol{Nois}$ datasets.
Methods MMLU PIQA SIQA GSM8K
HydraLoRA†(with HyNoIse) 40.57 73.40 45.03 8.47
HydraLoRA† 43.08 74.92 47.29 11.83
LoPE† 44.42 76.28 49.03 13.72
@@ 5.1 Does Higher HyNoise Ratio Enhance Performance?
In Section 4.2, while our baseline experiments employed a 5% HyNoIse ratio, we investigated whether an increased noise ratio during Stage I enhances noise robustness. We evaluated four HyNoIse ratios (3.5%, 5%, 10%, 15%), with comparative results visualized in Figure 4. Contrary to potential expectations, our analysis reveals a non-monotonic relationship between HyNoIse intensity and model robustness. Excessive noise injection (i.e. $\geq$ 10%) degrades the router’s discriminative capability for noise patterns, as evidenced by reduced expert specialization. We consider that this was due to overwhelming data distortion, which prevents effective differentiation between clean and noisy inputs, ultimately compromising expert weight allocation. We also evaluated the performance under different basic noise ratios; additional experimental results are provided in Appendix D.1.
@@ 5.2 Can the Poisoning Expert Truly Accomplish Its Task?
Method NE=3,PE=1 NE=3,PE=2 NE=2,PE=2 Mask(NE=3,PE=1) Not Mask(NE=3,PE=1)
@@ 6 Conclusion
This research is supported by the Sichuan International Science and Technology Innovation Cooperation Project (also known as the Hong Kong, Macao, and Taiwan Science and Technology Innovation Cooperation Project, ID: 2024YFHZ0317), the Chengdu Science and Technology Bureau Project (ID: 2024-YF09-00041-SN), the National Natural Science Foundation of China (Project ID: W2433163), Huawei Funding (Project ID: H04W241592), and the Postdoctoral Fellowship Program (Grade C) of the China Postdoctoral Science Foundation (Grant No. GZC20251053).